### 14 Mar 2024 v2

@@ Abstract
In many contexts of communication, the communication capabilities of the involved actors cannot be known beforehand, whether it is a cell, a plant, an insect, or even a life form unknown to Earth. Regardless of the recipient, the message space and time scale could be too fast, too slow, too large, or too small and may never be decoded. Therefore, it pays to devise a way to encode messages agnostic of space and time scales. We propose the use of fractal functions as self-executable infinite-frequency carriers for sending messages given their properties of structural self-similarity and scale invariance. We call it ‘fractal messaging’. Starting from a spatial embedding, we introduce a framework for a space-time scale-free messaging approach to this challenge. When considering a space and time-agnostic framework for message transmission, it would be interesting to encode a message such that it could be decoded at several spatio-temporal scales. Hence the core idea of the framework proposed herein is to encode a binary message as waves along infinitely many frequencies (in power-like distributions) and amplitudes, transmit such a message, and then decode and reproduce it. To do so, the components of the Weierstrass function, a known fractal, are used as carriers of the message. Each component will have its amplitude modulated to embed the binary stream, allowing for a space-time-agnostic approach to messaging.
@@ 2.1.2 The Weierstrass function
Figure 1 presents the general behaviour of $W_{a,b}(t)$ as the number of terms (or components) in the summation increase.
@@
Figure 1 indicates that $W_{a,b}(t)$ is nothing but the superposition of sinusoidal waves, which have been extensively used as carriers for messages using traditional radio transmission techniques. In the next section, a specific technique for amplitude modulation is discussed.
@@ 3 Intuition
Figure 3: Simple illustration of embedding $Re(z_{e})\geq 0$ (left) and $Im(z_{e})\geq 0$ (right) for $R_{e}=10$
@@ 4 Using the Weierstrass function’s components as carriers for binary messages
4.1 Example of encoded messages in fractal carriers
@@
Fractal behaviour is observed for the encoded message by zooming in the lower plot in Figure 4. These zoomed in regions are presented in Figure 5.
Figure 5: Zooming in $\mathcal{F}_{1/2,3}(^{\prime}hi!^{\prime},t)$. Twenty components were considered, for computational simplicity.
@@ 5 Conclusion
Our approach instantiates the idea of sending a self-encoding/self-decoding signal as a mathematical formula that carries both the data and its unfolding instructions. This would be similar to sending a self-executable message with a computer on which to run it. It is equivalent to sending self-executable computer code that, when unfolded, reads a message at all possible time scales and in all possible channels simultaneously.
One aspect to explore is how and whether we are prepared to receive and detect a signal with these properties. Together with our previous work [5], we have contributed to specific technical challenges at the receiver and now sender ends, and combining both approaches is to be done.

### 20 Mar 2024 v3

@@ Abstract
In many contexts of communication, the communication capabilities of the involved actors cannot be known beforehand, whether it is a cell, a plant, an insect, or even a life form unknown to Earth. Regardless of the recipient, the message space and time scale could be too fast, too slow, too large, or too small and may never be decoded. Therefore, it pays to devise a way to encode messages agnostic of space and time scales. We propose the use of fractal functions as self-executable infinite-frequency carriers for sending messages given their properties of structural self-similarity and scale invariance. We call it ‘fractal messaging’. Starting from a spatial embedding, we introduce a framework for a space-time scale-free messaging approach to this challenge. When considering a space and time-agnostic framework for message transmission, it would be interesting to encode a message such that it could be decoded at several spatio-temporal scales. Hence the core idea of the framework proposed herein is to encode a binary message as waves along infinitely many frequencies (in power-like distributions) and amplitudes, transmit such a message, and then decode and reproduce it. To do so, the components/cycles of the Weierstrass function, a known fractal, are used as carriers of the message. Each component will have its amplitude modulated to embed the binary stream, allowing for a space-time-agnostic approach to messaging.
@@ 2.1.2 The Weierstrass function
Figure 1 presents the general behaviour of $W_{a,b}(t)$ as the number of terms (or components) in the summation increase. The figure indicates that $W_{a,b}(t)$ is nothing but the superposition of sinusoidal waves, which have been extensively used as carriers for messages using traditional radio transmission techniques. In the next section, a specific technique for amplitude modulation is discussed.
@@ 3 Intuition
Figure 3: Simple illustration of embedding $Re(z_{e})\geq 0$ (left) and $Im(z_{e})\geq 0$ (right) for $R_{e}=10$, that is, the encoding of the message ‘01’ in the Mandelbrot set.
@@ 4 Using the Weierstrass function’s components as carriers for binary messages
4.1 Creating the fractal communication signal
In order to make full use of equation (3), the following pseudo-code can be used:
Algorithm 1 Generate $\mathcal{F}_{a,b}(msg,t)$
1:$a$, $b$, $M$ (number of components of $W_{a,b}(t)$ to be considered); $T$ (duration of the signal); $msg$ (message to be transmitted).
2:$msg_{binary}\Leftarrow$ a binary encoding for $msg$ (the ASCII binary representation of strings in $msg$, for example).
3:for $m\in\{0,1,...M\}$ do
4:     $C_{m}(t)\Leftarrow$ the $m$-th component of $W_{a,b}(t)$, which is a sinusoidal wave of amplitude $a^{m}$ and frequency $b^{m}/2$.
5:     $mask_{m}(t)\Leftarrow$ a mask composed of square waves juxtaposed (phase shifted) to represent each each bit in $msg_{binary}$. Each rectangular wave must have the same frequency as $C_{m}$.
6:     $masked_{m}(t)\Leftarrow C_{m}\times mask_{m}$, which represents the masked component.
7:end for
8:$\mathcal{F}_{a,b}(msg,t)\Leftarrow\sum_{m=0}^{M}masked_{m}(t)$.
9:Transmit $\mathcal{F}_{a,b}(msg,t)$ from $t=0$ to $t=T$.
@@ 4.1 Creating the fractal communication signal
In order to create the fractal message in Figure 4, Algorithm 1 was implemented in Python. Thus, to illustrate each step in the for loop, let one consider Figure 5.
Figure 5: First three components of the fractal message embedding the ASCII representation of “hi!”
Figure 5 presents the first three iterations of the for loop in Algorithm 1. For $C_{0}$, the 00-th component of the Weierstrass function, a simple sinusoidal wave of amplitude $(1/2)^{0}$ and frequency $23^{0}/$ 2 is seen. Then, the ASCII binary stream ‘011010000110100100100001’ is simply represented as a series of square waves side by side (one for each bit in the stream). This sequence of square waves is then multiplied by each element of the carrier (time-wise multiplication), leading to $masked_{0}$. For the $1$-th component, a similar process is carried out. Now, the amplitude of the sinusoidal wave is $(1/2)^{1}$ while its frequency is $23^{1}/$ 2. It can be seen that the square waves now have both their phase shifting and frequencies adjusted in order to match the properties of the $1$-th component. The same time-wise multiplication of elements is carried out to obtain $masked_{1}$. This process is also considered for the $2$-th component, with even smaller amplitudes and greater frequencies. After collecting $M$ components (defined according to the numerical precision sought), the complete fractal message can be obtained by summing all the $masked_{m}$, for $m\in\{0,1,...,M\}$. To produce the lower plot in Figure 4, 10 components were considered. Fractal behaviour is observed for the encoded message by zooming in the lower plot in Figure 4. These zoomed in regions are presented in Figure 6.
Figure 6: Zooming in $\mathcal{F}_{1/2,3}(^{\prime}hi!^{\prime},t)$. Twenty components were considered, for computational simplicity. They are all the same but at different scales, encoding all the same message ‘hi’ an infinite number of times.
@@ 5 Conclusion
Our approach instantiates the idea of sending a self-encoding/self-decoding signal as a mathematical formula that carries both the data and its unfolding instructions. This would be similar to sending a self-executable message with a computer on which to run it. It is equivalent to sending self-executable computer code that, when unfolded, reads a message at all possible time scales and in all possible channels simultaneously. Other fractal functions may be used and yet to explore is whether they offer any advantage given that the Weierstrass function seems to solve the main identified challenges. However, other fractals in higher dimensions may offer alternative embeddings for multi-modal non-linear messages.
Another aspect to explore is how and whether we are prepared to receive and detect a signal with these properties. Together with our previous work [5], we have contributed to specific technical challenges at the receiver and now sender ends, and combining both approaches is to be done.